• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:11:27 on August23,2003

    Entry number 112757

    keyword=Swing shift summary

    16:00-18:00 Beam tune and control access to replace the carbon target holders
    18:00-19:05 Beam orbit tunning
    19:20-19:30 Beam orbit tune (for Compton)
    19:45-20:05 NMR measurement =25% (the same as EPR)
    20:05 - RC: the shift plan is changed. We are going to take optics data and pump up the 3He cell polarization meanwhile

    20:35-21:00 moving to 2.813 GeV/c
    21:00-21:20 restarting the beam
    21:20-21:26 data taking
    21:26-21:50 moving to 2.704
    21:50-22:20 data taking
    22:30-23:05 moving to 2.992 GeV/c
    23:05 - NMR=28%
    23:10 - start data taking for kin 12.11
    23:40 - run 3919 3He 3M events - 1.8M "good" events, with polarizations of 35%x80% 6.6M "good" events are needed
    for one setting. With the current 28%x75.6% it is decided to take 80%
    of the equivalent statistics, namely 9.2M "good" events,
    which is 5 runs by 3M raw events.